Factor (x^2-14x+24)
Answer by babynarah2007(12)

Solution by EXPLAIN simplification of an expression
Your Result:
Simplify (x^2-14x+24)
you can solve by factor theory
(X-2)*(X-12)=X^2-2X-12X+24=X^2-12X+24 ( it is correct)

so, X=2 or X=12
very simply you have done!
Try your best